OpenStudy (anonymous):

I need to create a formula that represents if the food supply growth rate doubled using the following numbers --------------- In 2012, the population of Malthustan is 8 million people and increasing at a rate of 5% per year. Malthustan currently produces enough food to feed 12 million people, and is increasing its food supply to feed an additional 160,000 people per year. --- This is my guess, but I don't know if it's right... : y = 12000000 + 2(160000)x = 12000000 + 320000x

OpenStudy (anonymous):

Rules: 12,000,000 is currently being produced. 160,000 is added each year. Food Supply Growth Rate = (Annual Food Added to Supply)/(Current Food Produced) 1.) F.S.G.R. = 160,000/12,000,000 2.) 0.0134 = 160,000/12,000,000 Remember this is all based on the current food produced, so that will be our value for x. x= 12,000,000 Total Food Production will be: 1.) 0.0134(x) + x 2.) 0.0134(12,000,000) + 12,000,000 So, to double the current growth rate: 0.0134*2=0.0268 Our new formula is: 1.) 0.0268(x) + (x) 2.) 0.0268(12,000,000) + 12,000,000 To check our work: 1.) 160,000*2 = 320,000 2.) 0.0268(12,000,000) = 321,600 *off due to rounding of rate from .0133333333333 to .0134* I hope that helps.

OpenStudy (anonymous):

thanks for your response!!